A is now five times as old as B, but 7 year from now she will be 3 times, as old as he will be then how old is A now.

Answer by Cromlix(4381)

Hi there,
A is now 5 times as old as B
A = 5B
7 years from now she will be 3 times as old as B
A + 7 = 3(B + 7)
Replace A with 5B from 1st equation.
5B + 7 = 3B + 21
Collect like terms
5B - 3B = 21 - 7
2B = 14
B = 7
A is (7 x 5) = 35 yrs old now
Hope this helps:-)